# Online POMDP Planning with Anytime Deterministic Guarantees

### v1

@@ 5 Experiments
Figure 2: The graph illustrates the convergence of the deterministic bounds using Deterministically-Bounded-AR-DESPOT, with upper and lower bounds depicted alongside the optimal value obtained through exhaustive search.
In this section, we present the experimental results obtained by integrating deterministic bounds into a state-of-the-art algorithms namely, AR-DESPOT Somani et al. (2013) and POMCP Silver and Veness (2010) as a baseline. The primary objective of these experiments is to demonstrate the validity of our derived bounds, as presented in Theorem 2, and the corresponding algorithm outlined in Algorithm 1.
The implementation of our algorithm was carried out using the Julia programming language and evaluated through the Julia POMDPs package, provided by Egorov et al. (2017). Although the library primarily focused on infinite horizon problems, we made the required modifications to extend its capabilities to accommodate finite-horizon POMDPs. The experiments were conducted on a computing platform consisting of an Intel(R) Core(TM) i7-7700 processor with 8 CPUs operating at 3.60GHz and 15.6 GHz. The selection of hyper-parameters for the POMCP and AR-DESPOT solvers, and further details about the POMDPs used for our experiments are detailed in the appendix.
In Figure 2, we demonstrate the empirical convergence of the deterministic bounds in relation to planning time. For these experiments, we focused on a toy example, Tiger POMDP Kaelbling et al. (1998). By conducting an exhaustive search and computing a full posterior belief for each node, we derived the optimal value, depicted as a dashed green line in figure 2. The graph presents the convergence of the bounds calculated with Deterministically-Bounded AR-DESPOT (DB-DESPOT), to the optimal value. The mean values of the upper and lower bounds across 100 simulations are plotted, accompanied by the standard deviation, for time increments of $\Delta t=0.1$ seconds.
Algorithms that provide probabilistic guarantees have a non-zero probability of taking a suboptimal action regardless of the planning time. In contrast, when given sufficient planning time, the deterministic bounds can certify the optimal action once the lower bound of an immediate action exceeds the upper bounds of all other actions. In our experiments, we observed that although the baseline algorithms, AR-DESPOT and POMCP, and the deterministically bounded algorithms, DB-DESPOT and DB-POMCP, had access to the same tree and samples, AR-DESPOT and POMCP occasionally made incorrect decisions, resulting in degraded performance.
We evaluated the performance of both algorithms on different POMDPs, including the Tiger POMDP, Discrete Light Dark Sunberg and Kochenderfer (2018) and Baby POMDP. The corresponding results are summarized in Table 1. After each planning session the calculated best action is executed, the belief is updated according to the captured observation and a new planning session is invoked. The table reports the empirical mean and std of the cumulative rewards obtained by executing the calculated best actions based on $100$ runs.
Since the examined POMDPs have manageable sizes, DB-DESPOT could find the optimal action within the given time budget, limited to 1[sec]. An optimal action is determined when a lower bound of a given action is higher than any other upper bound; notably, this does not necessitates a construction of the entire planning tree and can be used as an early stopping mechanism for the optimal action. In the larger Laser Tag POMDP Somani et al. (2013), the DB-DESPOT did not outperform AR-DESPOT. This discrepancy occurred because the planning time was insufficient to guarantee an optimal action due to larger state, action and observation spaces of Laser Tag POMDP. While our proposed algorithms shows promise in more compact POMDPs, their performance in larger-scale problems, like the Laser Tag POMDP, especially when constrained by external time budgets, merits further investigation.
Table 1: Performance comparison with and without deterministic bounds.
Algorithm Tiger POMDP Laser Tag Discrete Light Dark Baby POMDP
[…]
6 Conclusions
In this work, we presented a novel methodology aimed at offering anytime, deterministic guarantees for approximate POMDP solvers. These solvers strategically leverage a subset of the state and observation spaces to alleviate the computational overhead. Our key proposition elucidates a linkage between the optimal value function, which is inherently computationally intensive, and a more tractable approximation frequently employed in contemporary algorithms. In the first part of the paper, we derived the theoretical relationship between the use of a selective subset of observations in a planning tree. One contribution of this work is the formulation of an upper deterministic bound (UDB) which governs exploration within the belief space, and is theoretically guaranteed to converge to the optimal value. This approach, however, depends on a complete belief update at each node of the tree, a requirement that can be computationally infeasible in many practical POMDPs. In the second part, we address this computational challenge, by extending our derivations to account for both a subset of the observations and states. This extension increases the feasibility of our approach and allows it to be incorporated into existing state-of-the-art algorithms. We have outlined how our methodology can be integrated within these algorithms. To illustrate the practical utility of our derivations, we applied them to certify and improve the solution of AR-DESPOT and POMCP algorithms, a state-of-the-art solvers for POMDPs.
6.1 Limitations and future work
While the application of our deterministic bounds within the context of the AR-DESPOT algorithm showcased its potential, there are several challenges that need to be addressed in future work. Firstly, the convergence rate of our proposed method remains an area for further exploration. A detailed theoretical analysis on this aspect could provide valuable insights into the practicality of our approach for larger, more complex domains. Secondly, the current use of the UDB is predominantly restricted to problems characterized by low state dimensionality. Extending the UDB to handle higher state dimensionality is a significant challenge due to the increased computational complexity involved. Finally, we consider the integration of the UDB with simplified state and observation spaces to be a promising future direction. This could potentially lead to a more comprehensive and efficient strategy for handling larger POMDPs, thereby further enhancing the applicability of our deterministic approach.

### v4

@@ 10 Experiments
10.1 POMDP scenarios
We begin with a brief description of the Partially Observable Markov Decision Process (POMDP) scenarios implemented for the experiments. each scenario was bounded by a finite number of time steps used for every episode, where each action taken by the agent led to a decrement in the number of time steps left. After the allowable time steps ended, the simulation was reset to its initial state.
10.1.1 Tiger POMDP
The Tiger is a classic POMDP problem Kaelbling et al. [1998], involves an agent making decisions between two doors, one concealing a tiger and the other a reward. The agent needs to choose among three actions, either open each one of the doors or listen to receive an observation about the tiger position. In our experiments, the POMDP was limited horizon of 5 steps. The problem consists of 3 actions, 2 observations and 2 states.
10.1.2 Discrete Light Dark
Is an adaptation from Sunberg and Kochenderfer [2018]. In this setting the agent needs to travel on a 1D grid to reach a target location. The grid is divided into a dark region, which offers noisy observations, and a light region, which offers accurate localization observations. The agent receives a penalty for every step and a reward for reaching the target location. The key challenge is to balance between information gathering by traveling towards the light area, and moving towards the goal region.
10.1.3 Laser Tag POMDP
In the Laser Tag problem, Somani et al. [2013], an agent has to navigate through a grid world, shoot and tag opponents by using a laser gun. The main goal is to tag as many opponents as possible within a given time frame. The grid is segmented into various sections that have varying visibility, characterized by obstacles that block the line of sight, and open areas. There are five possible actions, moving in four cardinal directions (North, South, East, West) and shooting the laser. The observation space cardinality is $\left|\mathcal{Z}\right|\approx 1.5\times 10^{6}$, which is described as a discretized normal distribution and reflect the distance measured by the laser. The states reflect the agent’s current position and the opponents’ positions. The agent receives a reward for tagging an opponent and a penalty for every movement, encouraging the agent to make strategic moves and shots.
10.1.4 Baby POMDP
The Baby POMDP is a classic problem that represents the scenario of a baby and a caregiver. The agent, playing the role of the caregiver, needs to infer the baby’s needs based on its state, which can be either crying or quiet. The states in this problem represent the baby’s needs, which could be hunger, discomfort or no need. The agent has three actions to choose from: feeding, changing the diaper, or doing nothing. The observations are binary, either the baby is crying or not. The crying observation does not uniquely identify the baby’s state, as the baby may cry due to hunger or discomfort, which makes this a partially observable problem. The agent receives a reward when it correctly addresses the baby’s needs and a penalty when the wrong action is taken.
10.2 Hyperparameters
The hyperparameters for both DB-DESPOT and AR-DESPOT algorithms were selected through a grid search. We explored an array of parameters for AR-DESPOT, choosing the highest-performing configuration. Specifically, the hyperparameter $K$ was varied across $\{10,50,500,5000\}$, while $\lambda$ was evaluated at $\{0,0.01,0.1\}$. Similarly, DB-POMCP and POMCP were examined three different values for the exploration-exploitation weight, $c=\{0.1,1.0,10.0\}$ multiplied by $V_{max}$, which denotes an upper bound for the value function.
For the initialization of the upper and lower bounds used by the algorithms, we used the maximal reward, multiplied by the remaining time steps of the episode, $\mathcal{R}_{\max}\cdot(\mathcal{T}-t-1)$.